Circular No. 7577

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2000 Y10 (MUELLER 4) S. Nakano, Sumoto, Japan, reports the recovery of P/1992 G3 (= 1992g

= 1992 IV) by T. Oribe (Saji Observatory) on CCD images taken with a 1.03-m reflector, with the astrometry given below. The comet is faint and of stellar appearance on 2000 Dec. 22 (m_{2} = 20.5). The indicated correction to the orbital elements on MPC 31663 (ephemeris on MPC 41213) is ΔT = +0.23 day.

2000/01 UT  α_{2000}  δ_{2000}  m_{1}
Dec. 22.84689  13^{h}16^{m}40.^{s}36  + 7^{o}38′25′′.1
22.85209  13 16 40.80  + 7 38 26.1
22.85765  13 16 41.34  + 7 38 27.5
22.86251  13 16 41.83  + 7 38 27.0
Jan. 30.83403  14 06 40.26  +11 45 20.1  19.6
30.83750  14 06 40.45  +11 45 22.0
30.84167  14 06 40.68  +11 45 24.9

Nakano provides the following orbital elements from 35 observations, 1992–2001 (mean residual 0′′.8):

Epoch = 2001 Feb. 20.0 TT

T = 2001 Feb. 7.8847 TT  ω = 43.^{o}6165
e = 0.388590  Ω = 145.3691 }2000.0
q = 2.646731 AU  i = 29.7486

a = 4.328900 AU n^{o} = 0.1094304 P = 9.007 years

SUPERNOVA 2001B IN IC 391 R. Chornock and A. V. Filippenko, University of California at Berkeley,

report that inspection of fully calibrated CCD spectra (range 330–1000 nm) of SN 2001B (cf. IAUC 7555), obtained on Jan. 23 UT with the Shane 3-m reflector at Lick Observatory, reveals that the object is probably a type-Ib supernova, roughly one week past maximum brightness (not type-Ia, as reported on IAUC 7563). The spectrum exhibits well-developed absorption lines of He I, as well as an unidentified absorption line near 625 nm, similar to one seen in SN 2000dt (cf. IAUC 7511).

2001 February 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green